OpenStudy (amistre64):

to find the range of f(x,y), would we find that partials fx and fy equal to 0 ?

OpenStudy (anonymous):

I think that would give you the locations where the rate of change is 0 with respect to each variable rather than the range.

OpenStudy (amistre64):

right, it could be a saddleback; a min or a max ....

OpenStudy (amistre64):

fxx and fyy would determine cave i think

OpenStudy (anonymous):

Ah, maybe I'm misunderstanding range. What does it mean in this context?

OpenStudy (amistre64):

it would provide a point of interest; the actual value would depend on plugging x and y back into f(x,y)

OpenStudy (amistre64):

all possible outputs ...

OpenStudy (anonymous):

That would be a way to find those points of interest, so it would give you the range of the function if you had a "nice" function.

OpenStudy (amistre64):

spose f(x,y) = ln(4-x^2-9y^2) determine domain and range

OpenStudy (amistre64):

nice functions dont exist lol

OpenStudy (anonymous):

Lol, I meant differentiable.

OpenStudy (amistre64):

fx = -2x/(4-x^2-9y^2) fy = -18y/(4-x^2-9y^2) at x=0 and y=0 we have a max since this things an eggy ellipse range is up to ln(4)

OpenStudy (amistre64):

but is the lower limit -inf? or 0?

OpenStudy (anonymous):

That makes sense to me, but I'm not sure how to find out that lower limit.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=range+f%28x%2Cy%29%3Dln%284-x%5E2-9y%5E2%29 the wolf says unable to determine range lol

OpenStudy (anonymous):

Yeah, I looked at that, too, lol.

OpenStudy (anonymous):

Looking at it intuitively, 4-x^2-9y^2 can get arbitrarily close to 0, which should mean that ln(4-x^2-9y^2) should be able to approach negative infinity.

OpenStudy (amistre64):

which is what I was thinking as well; which is why there isnt another fx or fy to zero out of

OpenStudy (amistre64):

http://web.monroecc.edu/manila/webfiles/calcNSF/JavaCode/CalcPlot3D.htm thats useful

OpenStudy (anonymous):

This is sexy.

OpenStudy (amistre64):

lol